# Secret-Key Generation from Private Identifiers under Channel Uncertainty

### v1

@@ Abstract
Index Terms:
Capacity region, compound channels, multiple outputs, key generation, privacy leakage, PUFs.
@@ I Introduction
Secret-key generation using physical identifiers, such as responses from unclonable functions (PUFs), is a promising protocol for device authentication because it offers several advantages, including simple designs, low costs, and eliminating the need to save the secret key on the device [3]. A PUF is defined as a physical function that for a given input (challenge), provides an output (response) that serves as a unique identifier for each device [4, 5]. Some examples of PUFs are static random access memory (SRAM) PUFs and ring-oscillator (RO) PUFs.
@@
In subsequent discussions, the response of a PUF unaffected by noise is referred to as the source identifier. The responses of a PUF observed at the terminals and Eve through communication channels are called the observed identifiers.
@@ I-B Related Work
Secret-key generation using PUFs [^2] PUF and biometric identifiers share similar characteristics, and thus, the theoretical results developed for one can be applied to the other as well [3]. has been studied from information-theoretic perspectives in [8, 9]. Later, several extensions of this model were found in [10, Ch. 4] for Gaussian sources, [11] for separated and combined enrollments, and [12] for multiple rounds of enrollments and authentications. Limited storage rate was introduced to the model in [13]. Furthermore, the fundamental limits among secret-key, storage, and privacy-leakage rates when Eve also has a correlated sequence of the source were characterized in [14] for discrete sources and [15] for Gaussian sources. Note that this model is similar to the key-generation problem [16], [17], [18], but an additional privacy constraint is imposed in the problem formulation to limit information leakage on the source identifier.
@@
Secret-key generation with PUFs for compound sources has been studied in [26, 27] for the generated-secret (GS) model and the chosen-secret (CS) model [8, Sect. III]. Some extensions on this setting are explored in [28, 29] to incorporate user identification. Similar problems can be found in [30, 31, 32, 33] for key generation where the privacy constraint is not imposed and in [34, 35, 36, 37] for compound wiretap channels.
@@ I-C Main Challenges and Contributions
The setup in [26, 27] focuses on uncertainty of the source statistics, whereas our work addresses channel uncertainty, when the channels can have multiple states whose realizations are unknown to the terminals. Additionally, compared to these prior studies [26, 27], our model incorporates enrollment noise and side information at Eve. Moreover, while previous works [26, 27] strictly focus on discrete sources, our study extends the discussion to investigate the capacity regions for both discrete and Gaussian cases. Below, we highlight some technical challenges encountered during the analysis of our newly considered setting.
For general discrete sources, since the encoder and decoder do not know the true CSI, in the achievability proof, the secret key’s uniformity and secrecy-leakage constraints must be satisfied for all possible receiver-eavesdropper channel states. In point-to-point communications [20], these constraints can be handled by bounding conditional entropy for joint typical sequences. However, in our analysis, an additional step, showing that there exists a random variable that simultaneously achieves the maximum mutual information across all Eve’s channels and is jointly typical with the source random variables with high probability, is essential to analyze these constraints. Moreover, unlike [20], the obtained inner and outer bounds involve optimizations over the channel states. Specifically, the inner region first involves an optimization carried over the distributions of auxiliary random variables, and then a minimization of the index pair for channel states. In contrast, the order of these two operations is reversed in the outer region. This leads to a gap between the inner and outer bounds, similar to the conclusion drawn in [34] for compound wiretap channels.
@@
However, we show that, for a noiseless enrollment channel, our inner and outer regions coincide for Gaussian sources, providing a complete capacity characterization. The main challenging aspect arises in proving the converse part. Given the multiple-antenna settings at the legitimate terminals and Eve, the vector-form observations are not stochastically degraded in general [38]. We use sufficient statistics [39] to convert the vector problem into a scalar one. However, after this conversion, showing that all constraints of the original problem definition are preserved is challenging, and it is unclear whether the same expressions of the outer bounds for general discrete sources also hold for the scalar variables. Therefore, we cannot directly apply the technique in [40, Appx. B] to eliminate the second auxiliary random variable. In this paper, we instead derive new single-letter expressions of outer bounds for the Gaussian case using scalar random variables.
@@
Another difficulty arises in proving the converse for the parametric expression of Gaussian sources. In the analysis of the model without side information at Eve [10, Appx. D], [41], the conditional entropy power inequality (EPI) plays an important role. However, the EPI is insufficient to prove the converse for all possible values of the optimization parameter in our problem. To overcome this issue, we adopt a distinct method introduced in [42, Sect. IV-C], using Fisher information. This approach enables us to derive the outer region that coincides with the inner one for any value of the optimization parameter.
@@
  • We provide a complete characterization of the capacity regions of the GS and CS models for Gaussian sources by demonstrating the existence of a saddle point at which the inner and outer bounds coincide. Moreover, numerical calculations for the GS model are provided to illustrate how the change of the number of antennas at the decoder and Eve affects the secret-key and storage rates.
@@
Our results recover, as special cases, results derived in previous works. For discrete sources, when only the channel to the decoder is compound and all channels have a single output, the inner and outer bounds match the preliminary result given in [1, Props. 1 and 2]. Additionally, the inner and outer bounds are tight for single-output and non-compound channels, and recover [20, Th. 3 and 4] without action cost. For Gaussian sources, as detailed in Section III-B, our results recover as special cases the capacity regions derived in [10, Ch. 4] and [21] for single-output and non-compound channels.
I-D Notation and Paper Organization
@@
$\mathbb{R}_{+}$ is the set of non-negative real numbers. For any $a,b\in\mathbb{R}$, define $[a:b]\triangleq[\lfloor a\rfloor,\lceil b\rceil]\cap\mathbb{N}$. Italic uppercase $X$ and lowercase $x$ denote a random variable and its realization, respectively. Boldface letters $\bf X$ and $\bf x$ represent a collection of random variables and its realization. $X^{n}$ denotes the vector $(X_{1},{\dots},X_{n})$ and $X_{t}$ represents the $t$-th element in the vector. $X_{k}^{t}$ stands for a partial sequence $(X_{k},\dots,X_{t})$ for any $[k:t]\subseteq[1:n]$. $\sigma^{2}_{X}$ and ${\bf\Sigma}_{\mathbf{Y}}$ denote the variance of $X$ and the covariance matrix of ${\bf Y}$. $\mathcal{N}(0,\sigma^{2})$ denotes the Gaussian distribution with zero mean and variance $\sigma^{2}$. $\mathcal{T}^{n}_{\delta}(X)$ denotes the set of $\delta$-strongly typical sequences according to $P_{X}$ [43] and the random variable inside the parentheses is omitted, e.g., $\mathcal{T}^{n}_{\delta}$, when it is clear from the context. Additionally, the set of conditionally $\delta$-typical sequences is denoted as $\mathcal{T}^{n}_{\delta}(XY|z^{n})\triangleq\{(x^{n},y^{n}):(x^{n},y^{n},z^{n})\in\mathcal{T}^{n}_{\delta}\}$.
@@
The remainder of the paper is organized as follows. In Section II, we state the problem definitions for the GS and CS models. We present our main results in Section III. The proofs of our main results are available in the appendices. Finally, we give concluding remarks in Section IV.
@@ II Problem Statement
$\mathcal{R}_{G}$ is defined as the closure of the set of all achievable rate tuples for the GS model, and it is called the capacity region.
@@
A tuple of secret-key, storage, and privacy-leakage rates $(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}$ is achievable for the CS model if, for sufficiently small $\delta>0$ and large enough $n$, there exist pairs of encoders and decoders that satisfy all the conditions (2)–(6) with replacing (3) with $\log|\mathcal{S}|\geq n(R_{S}-\delta).$ Let $\mathcal{R}_{C}$ be the capacity region of the CS model.
@@ III Main Results
In this section, we first present inner and outer bounds for the capacity region of the GS and CS models with discrete sources in Section III-A. We then present the capacity regions for Gaussian sources in Section III-B.
@@ III-A Discrete Sources
$\mathcal{R}_{G}$ $\supseteq\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}}}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$
@@
$\mathcal{R}_{C}$ $\supseteq\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}}}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$
@@
where auxiliary random variables $V$ and $U$ satisfy the Markov chain $V-U-\tilde{\mathbf{X}}-X-(\mathbf{Y}_{k},\mathbf{Z}_{l})$ for all $k\in\mathcal{K}$ and $l\in\mathcal{L}$, and $|\mathcal{V}|\leq|\tilde{\mathbfcal{X}}|+6$ and $|\mathcal{U}|\leq(|\tilde{\mathbfcal{X}}|+6)(|\tilde{\mathbfcal{X}}|+3)$.
@@
See Appendix A-A. ∎
@@
$\mathcal{R}_{G}\subseteq$ $\bigcap_{k\in\mathcal{K}}\bigcap_{l\in\mathcal{L}}\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}}}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:~{}$
@@
$\mathcal{R}_{C}\subseteq$ $\bigcap_{k\in\mathcal{K}}\bigcap_{l\in\mathcal{L}}\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}{}}}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:~{}$
@@
where $V$ and $U$ satisfy the Markov chain $V-U-\tilde{\mathbf{X}}-X-(\mathbf{Y}_{k},\mathbf{Z}_{l})$, and $|\mathcal{V}|\leq|\tilde{\mathbfcal{X}}|+5$, $|\mathcal{U}|\leq(|\tilde{\mathbfcal{X}}|+5)(|\tilde{\mathbfcal{X}}|+3)$.
@@
See Appendix A-B. ∎
The region in (9) matches [1, Prop. 2] when $\Omega_{\tilde{X}}=\Omega_{Y}=\Omega_{Z}=|\mathcal{L}|=1$. Moreover, in the non-compound settings, i.e., when $|\mathcal{K}|=1=|\mathcal{L}|$, the bounds in Propositions 1 and 2 match, yielding the following result.
@@
$\mathcal{R}_{G}=$ $\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}}}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:~{}$
@@
$\mathcal{R}_{C}=$ $\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}{}}}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:~{}$
@@
$R_{L}$ $\geq I(X;U|\mathbf{Y})+I(\mathbf{Y};V)-I(\mathbf{Z};V)\Big{\}}.$ (12)
@@
=$\overset{\rm(b)}{=}I(X;U|\mathbf{Y})+I(\mathbf{Y};V)-I(\mathbf{Z};V),$ (15)
@@
The privacy-leakage rate in [20, Th. 3 and 4] without action cost is bounded as
@@
$=I(X;U|Y)+I(Y;V)-I(Z;V)+I(X;Z),$
where the equality holds by the Markov chain $V-X-(Y,Z)$. Compared to the privacy-leakage rate in (11) and (12), there is an extra term $I(X;Z)$, because the privacy-leakage rate constraint in [20, Th. 3 and 4] is defined as $\frac{1}{n}I(X^{n};J,Z^{n})=\frac{1}{n}I(X^{n};J|Z^{n})+I(X;Z)$. Therefore, (11) and (12) coincide with [20, Th. 3 and 4] (without action cost) if [20, eq. (5)] is replaced by (6).
In general, the regions in Propositions 1 and 2 do not coincide. More specifically, the inner bounds are obtained by optimizing the test channels $P_{V|U}$ and $P_{U|\tilde{\mathbf{X}}}$ for all possible channel states, while for the outer bounds, the intersection over the indices $k$ and $l$ is taken before the union of the test channels. Optimizing for an arbitrary state index pair could yield a larger region compared to the one for all possible channel pairs, leading to a gap between the inner and outer bounds, similar to the observation drawn for compound-wiretap channels in [34].

### v2

@@ I Introduction
Secret-key generation using physical identifiers, such as responses from physical unclonable functions (PUFs), is a promising protocol for device authentication because it offers several advantages, including simple designs, low costs, and eliminating the need to save the secret key on the device [3]. A PUF is defined as a physical function that for a given input (challenge), provides an output (response) that serves as a unique identifier for each device [4, 5]. Some examples of PUFs are static random access memory (SRAM) PUFs and ring-oscillator (RO) PUFs.
@@
In subsequent discussions, the response of a PUF unaffected by noise is referred to as the source identifier. The responses of a PUF observed at the terminals and Eve through communication channels are called the observed identifiers. It is worth noting that the source identifiers are assumed to be fixed and thus the secret-key generation model considered in this paper corresponds to the source-type model, unlike the channel-type models, where the source distribution can be controlled [7].
@@ I-B Related Work
Secret-key generation using PUFs [^2] PUF and biometric identifiers share similar characteristics, and thus, the theoretical results developed for one can be applied to the other as well [3]. has been studied from information-theoretic perspectives in [9, 10]. Later, several extensions of this model were found in [11, Ch. 4] for Gaussian sources, [12] for separated and combined enrollments, and [13] for multiple rounds of enrollments and authentications. Limited storage rate was introduced to the model in [14]. Furthermore, the fundamental limits among secret-key, storage, and privacy-leakage rates when Eve also has a correlated sequence of the source were characterized in [15] for discrete sources and [16] for Gaussian sources. This model is similar to the key-agreement problem with forward communication only studied in [17], [18], [19], but an additional privacy constraint is imposed in the problem formulation to limit information leakage on the source identifier.
@@
Secret-key generation with PUFs for compound sources has been studied in [27, 28] for the generated-secret (GS) model and the chosen-secret (CS) model. In the GS model, the secret key is generated using the observed identifier at the encoder. In contrast, the CS model assumes that the secret key is independently and uniformly chosen in advance. Relevant applications of the GS model include field-programmable gate array (FPGA) based key generation with PUFs [29, 30] and that of the CS model can be seen in key-binding biometric authentication [31] and fuzzy commitment schemes [32, 33]. Some extensions on this setting are explored in [34, 35] to incorporate user identification. Similar problems can be found in [36, 37, 38, 39] for key generation where the privacy constraint is not imposed and in [40, 41, 42, 43] for compound wiretap channels. We note that the works [40, 41, 42, 43] focus on compound channels under the channel-type model, whereas our work addresses compound structure in the source-type model.
@@ I-C Main Challenges and Contributions
We begin by explaining challenges of proving the achievability part for general discrete sources. In [37], key generation for compound sources without the privacy constraint is investigated. While [37] derives a single-letter inner bound for discrete sources and a single-letter outer bound for degraded sources, we establish single-letter inner and outer bounds for discrete sources and also characterize the capacity region for Gaussian sources, which require different approaches from the discrete case. The key differences for inner-bound derivations between the work [37] and ours are twofold.
First, the techniques used for analyzing the secret-key uniformity and secrecy-leakage constraints are distinct. In [37, Th. 1], the secret key is derived from the shared randomness between encoder and decoder, and the analyses of the two constraints rely on extending the method proposed in [44] for non-compound sources. Our approach, in contrast, aligns with [21], where the secret key is generated through index mapping, and the analyses of the constraints build upon the technique used in [40] for analyzing the secrecy constraint in compound wiretap channels.
Second, the privacy-leakage constraint is not considered in [37]. In our problem formulation, as in [21, 22], this constraint is imposed and quantified by the mutual information between the source identifier and the helper data, conditioned on Eve’s observation. Its analysis is not straightforward because the helper data does not have an independent and identically distributed (i.i.d.) structure: although the encoder observes an i.i.d. identifier sequence, it generates the helper data based on the entire block rather than on individual symbols.
In the converse part, for a given channel state, the proofs of the GS and CS models mirror the ones in [21, Th. 3 and 4] with a proper modification for the privacy-leakage analysis as the definition is distinct. These results are then generalized to the compound-channel settings by taking the intersection over all possible channel states to establish the outer bounds. As a result, the inner region first involves an optimization carried over the distributions of auxiliary random variables, and then a minimization of the index pair for channel states. In contrast, the order of these two operations is reversed in the outer region. This leads to a gap between the inner and outer bounds, similar to the conclusion drawn in [40] for compound wiretap channels.
@@
However, we show that, for a noiseless enrollment channel, our inner and outer regions coincide for Gaussian sources, providing a complete capacity characterization. The main challenging aspect arises in proving the converse part. Given the multiple-antenna settings at the legitimate terminals and Eve, the vector-form observations are not stochastically degraded in general [45]. We use sufficient statistics [46] to convert the vector problem into a scalar one. However, after this conversion, showing that all constraints of the original problem definition are preserved is challenging, and it is unclear whether the same expressions of the outer bounds for general discrete sources also hold for the scalar variables. Therefore, we cannot directly apply the technique in [47, Appx. B] to eliminate the second auxiliary random variable. In this paper, we instead derive new single-letter expressions of outer bounds for the Gaussian case using scalar random variables.
@@
Another difficulty arises in proving the converse for the parametric expression of Gaussian sources. In the analysis of the model without side information at Eve [11, Appx. D], [48], the conditional entropy power inequality (EPI) plays an important role. However, the EPI is insufficient to prove the converse for all possible values of the optimization parameter in our problem. To overcome this issue, we adopt a distinct method introduced in [49, Sect. IV-C], using Fisher information. This approach enables us to derive the outer region that coincides with the inner one for any value of the optimization parameter.
@@
  • We provide complete characterizations of the capacity regions of the GS and CS models for Gaussian sources by demonstrating the existence of a saddle point at which the inner and outer bounds coincide.
  • We conduct numerical calculations for the Gaussian case to illustrate how the change of the number of antennas at the decoder and Eve affects the secret-key and storage rates. The results show that increasing the number of antennas at the decoder leads to a higher secret-key rate, while increasing antennas at Eve reduces the secret-key rate. Nevertheless, even if Eve has more antennas, a positive secret-key rate is still achievable as long as the worst channel power gain at the decoder is greater than the best channel power gain at Eve. Moreover, we compare the secret-key and privacy-leakage rates between the GS and CS models under the same values of storage rates. The results reveal that in the low storage-rate regime, the GS model outperforms the CS model in terms of secret-key rate, whereas the CS model provides better privacy-leakage performance. In the high storage-rate regime, the CS model is better suited as it can achieve the same secret-key rate as the GS model but with lower privacy leakage.
@@
Our results recover, as special cases, results derived in previous works. For discrete sources, when only the channel to the decoder is compound and all channels have a single output, the inner and outer bounds match the preliminary result given in [1, Props. 1 and 2]. Additionally, the inner and outer bounds are tight for single-output and non-compound channels, and recover [21, Th. 3 and 4] without action cost. For Gaussian sources, as detailed in Section III-B, our results recover as special cases the capacity regions derived in [11, Ch. 4] and [22] for single-output and non-compound channels.
I-D Modeling Assumptions
In general, PUF responses from devices are correlated. However, techniques such as transform coding-based algorithms [3] and principal component analysis [50] can be applied to convert these responses into a sequence with almost independent symbols. Therefore, we assume that each symbol in the source and observed identifier sequences is i.i.d. generated. Additionally, we assume that the database that stores helper data is public, e.g., in the cloud, and accessible to both the decoder and Eve. These modeling assumptions are consistent with those used in prior works [9, 10, 20, 21, 22].
I-E Notation and Paper Organization
@@
$\mathbb{R}_{+}$ is the set of non-negative real numbers. For any $a,b\in\mathbb{R}$, define $a:b]\triangleq[\lfloor a\rfloor,\lceil b\rceil]\cap\mathbb{N}$. Italic uppercase $X$ and lowercase $x$ denote a random variable and its realization, respectively. Boldface letters $\bf X$ and $\bf x$ represent a collection of random variables and its realization. $X^{n}$ denotes the vector $(X_{1},{\dots},X_{n})$ and $X_{t}$ represents the $t$-th element in the vector. $X_{k}^{t}$ stands for a partial sequence $(X_{k},\dots,X_{t})$ for any $k:t]\subseteq[1:n]$. $\sigma^{2}_{X}$ and ${\bf\Sigma}_{\mathbf{Y}}$ denote the variance of $X$ and the covariance matrix of $\bf Y$}. $\mathcal{N}(0,\sigma^{2})$ denotes the Gaussian distribution with zero mean and variance $\sigma^{2}$. $\mathcal{T}^{n}_{\delta}(X)$ denotes the set of $\delta$-strongly typical sequences according to $P_{X}$ [51] and the random variable inside the parentheses is omitted, e.g., $\mathcal{T}^{n}_{\delta}$, when it is clear from the context. Additionally, the set of conditionally $\delta$-typical sequences is denoted as $\mathcal{T}^{n}_{\delta}(XY|z^{n})\triangleq\{(x^{n},y^{n}):(x^{n},y^{n},z^{n})\in\mathcal{T}^{n}_{\delta}\}$ for a given $z^{n}\in\mathcal{Z}^{n}$.
@@
The remainder of the paper is organized as follows. In Section II, we state the problem definitions for the GS and CS models. We present our main results in Section III. Proofs of our main results are available in the appendices. Finally, we provide concluding remarks and some future directions in Section IV.
@@ II Problem Statement
$\mathcal{R}_{\rm G}$ is defined as the closure of the set of all achievable rate tuples for the GS model, and it is called the capacity region.
@@
A tuple of secret-key, storage, and privacy-leakage rates $(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}$ is achievable for the CS model if, for sufficiently small $\delta>0$ and large enough $n$, there exist pairs of encoders and decoders that satisfy all the conditions (2)–(6) with replacing (3) by $\log|\mathcal{S}|\geq n(R_{S}-\delta).$ Let $\mathcal{R}_{\rm C}$ be the capacity region of the CS model.
@@ III Main Results
This section presents the inner and outer bounds for the GS and CS models with discrete sources, followed by the tight bounds for Gaussian sources and numerical results for both models.
@@ III-A Discrete Sources
$\mathcal{R}_{\rm G}$ $\supseteq\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}}}\Big\{(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$
@@
$\mathcal{R}_{\rm C}$ $\supseteq\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}}}\Big\{(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$
@@
where auxiliary random variables $V$ and $U$ satisfy the Markov chain $V-U-\tilde{\mathbf{X}}-X-(\mathbf{Y}_{k},\mathbf{Z}_{l})$ for all $k\in\mathcal{K}$ and $l\in\mathcal{L}$, and $|\mathcal{V}|\leq|\tilde{\mathbfcal{X}}|+6$ and $|\mathcal{U}|\leq(|\tilde{\mathbfcal{X}}|+2(|\mathcal{K}|+|\mathcal{L}|)+1)(|\tilde{\mathbfcal{X}}|+|\mathcal{K}|+|\mathcal{L}|)+1)$.
@@
The proof is available in Appendix A-A, where the random codebook is constructed based on two layered random coding techniques. The first layer consists of the auxiliary sequences $V^{n}$ generated by $P_{V}$ and the second layer consists of the auxiliary sequences $U^{n}$ associated with $P_{U|V}$. The main challenge in the proof is to ensure that the secret-key uniformity (3) and the secrecy-leakage constraint (5) are satisfied for all possible receiver-eavesdropper states. To prove these constraints, the key idea is to introduce a random variable $\tilde{\mathbf{Z}}^{n}$ that jointly satisfies the equality $I(\tilde{\mathbf{Z}};U|V)=\max_{l}I(\mathbf{Z}_{l};U|V)$ and the Markov chain $V-U-\tilde{\mathbf{X}}-\tilde{\mathbf{Z}}$. The random variable $\tilde{\mathbf{Z}}^{n}$ plays a central role in analyzing the two constraints. This technique is not seen in the existing works [9, 21, 22] that study the secret-key generation with PUFs without compound channels. ∎
In Proposition 1, how each term in the constraints defining the regions $\mathcal{R}_{\rm G}$ and $\mathcal{R}_{\rm C}$ arises can be explained as follows. We begin with the region $\mathcal{R}_{\rm G}$. In the secret-key rate constraint, the term $\min_{k}I(\mathbf{Y}_{k};U|V)$ represents the minimum rate required for reliably estimating the sequence $U^{n}$ across all indices $k$, which in turn enables reliable reconstruction of the secret key since the key is extracted from $U^{n}$. On the other hand, the term $\max_{l}I(\mathbf{Z}_{l};U|V)$ is the maximum rate at which Eve can gain information about $U^{n}$ over all indices $l$. Therefore, the achievable secret-key rate is given by the difference $\min_{k}I(\mathbf{Y}_{k};U|V)-\max_{l}I(\mathbf{Z}_{l};U|V)$, similar to the one derived in [37, Th. 1] for compound sources. The terms $\max_{k}I(\tilde{\mathbf{X}};V|\mathbf{Y}_{k})$ and $\max_{k}I(\tilde{\mathbf{X}};U|V,\mathbf{Y}_{k})$ in the storage-rate constraint represents the rates of the bin indices at the first and second layers, respectively. In each layer, the maximum rate across all indices $k$ must be shared between the encoder and decoder to ensure reliable reconstruction of the secret key at the decoder. For the privacy-leakage rate, note that we can expand the mutual information $\frac{1}{n}\max_{l}I(X^{n};J|\mathbf{Z}^{n}_{l})$ as $\frac{1}{n}H(J)-\frac{1}{n}H(J|X^{n})-\frac{1}{n}\min_{l}I(\mathbf{Z}^{n}_{l};J)$ by using the Markov chain $J-X^{n}-\mathbf{Z}^{n}_{l}$. In the constraint of the privacy-leakage rate, the first and second terms in the right-hand side represent the upper bound on the entropy $\frac{1}{n}H(J)$, the third term represents the upper bound on the conditional entropy $\frac{1}{n}H(J|X^{n})$, and the forth and fifth terms represent the upper bound on the mutual information $\frac{1}{n}\min_{l}I(\mathbf{Z}^{n}_{l};J)$.
For the region $\mathcal{R}_{\rm C}$, the codebook and coding scheme developed for proving the region $\mathcal{R}_{\rm G}$ are employed as a subsystem to prove the achievability part. One-time pad operation is applied to conceal the chosen secret key in the CS model by adding the secret key generated in the subsystem [9, Appx. B-C], which leads to the same achievable secret-key rate. However, the storage rate is different because the masked information must be saved in the public database together with the helper data generated by the subsystem, so that the chosen secret key can be reliably estimated at the decoder. Therefore, the storage rate of the CS model is the sum of the storage rate for the GS model (the subsystem) and the secret-key rate. Moreover, the privacy-leakage rate remains unchanged because the concealed information reveals no extra leakage to Eve after applying the one-time pad addition. Similar behaviors are reflected in the outer bound derived below in Proposition 2.
@@
$\mathcal{R}_{\rm G}\subseteq$ $\bigcap_{k\in\mathcal{K}}\bigcap_{l\in\mathcal{L}}\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}}}\Big\{(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$~
@@
$\mathcal{R}_{\rm C}\subseteq$ $\bigcap_{k\in\mathcal{K}}\bigcap_{l\in\mathcal{L}}\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}{}}}\Big\{(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$~
@@
where $V$ and $U$ satisfy the Markov chain $V-U-\tilde{\mathbf{X}}-X-(\mathbf{Y}_{k},\mathbf{Z}_{l})$, and $|\mathcal{V}|\leq|\tilde{\mathbfcal{X}}|+2(|\mathcal{K}|+|\mathcal{L}|)+1$, $|\mathcal{U}|\leq(|\tilde{\mathbfcal{X}}|+2(|\mathcal{K}|+|\mathcal{L}|)+1)(|\tilde{\mathbfcal{X}}|+|\mathcal{K}|+|\mathcal{L}|+1)$.
@@
The proof is provided in Appendix A-B. For a fixed state $(k,l)$, the proof is the same as that of [21, Th. 3 and 4] without considering the action cost. Therefore, we make use of the result of those theorems to derive Proposition 2 for the compound channel setting. However, due to the difference in the definition of the privacy-leakage rate, appropriate modifications are required. ∎
The region in (9) matches [1, Prop. 2] when $\Omega_{\tilde{X}}=\Omega_{Y}=\Omega_{Z}=|\mathcal{L}|=1$. Moreover, in the non-compound settings, i.e., when $|\mathcal{K}|=1=|\mathcal{L}|$, the bounds in Propositions 1 and 2 match, yielding the following corollary.
@@
$\mathcal{R}_{\rm G}=$ $\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}}}\Big\{(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$~
@@
$\mathcal{R}_{\rm C}=$ $\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}{}}}\Big\{(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$~
@@
$R_{L}$ $\geq I(X;U|\mathbf{Y})+I(\mathbf{Y};V)-I(\mathbf{Z};V)\Big\},$ (12)
@@
$\overset{\rm(b)}{=}I(X;U|\mathbf{Y})+I(\mathbf{Y};V)-I(\mathbf{Z};V),$ (15)
@@
The privacy-leakage rate in [21, Th. 3 and 4] without action cost is bounded as
@@
$=I(X;U|Y)+I(Y;V)-I(Z;V)+I(X;Z),$ (16)
where the equality holds by the Markov chain VX(Y,Z)V-X-(Y,Z). Compared to the privacy-leakage rate in (11) and (12), there is an extra term $I(X;Z)$, because the privacy-leakage rate constraint in [21, Th. 3 and 4] is defined as $\frac{1}{n}I(X^{n};J,Z^{n})=\frac{1}{n}I(X^{n};J|Z^{n})+I(X;Z)$. Therefore, (11) and (12) coincide with [21, Th. 3 and 4] (without action cost) if [21, eq. (5)] is replaced by (6).
In Propositions 1 and 2, the orders of the optimization (union) over the test channels $P_{V|U},P_{U|\tilde{\mathbf{X}}}$ and the minimization (intersection) over the channel states $(k,l)$ are reversed. Specifically, Proposition 1 requires one to choose test channels $P_{V|U},P_{U|\tilde{\mathbf{X}}}$ that work simultaneously for all $(k,l)$ pairs, whereas Proposition 2 allows one to choose different test channels $P_{V|U},P_{U|\tilde{\mathbf{X}}}$ for a channel state pair $(k,l)$, and then only keeps the intersection over what is achievable per channel state pair. Therefore, Proposition 1 imposes stronger requirements, and as a result, the regions in Proposition 2 may be potentially larger.